### v1

# PuzzleWorld: A Benchmark for Multimodal, Open-Ended Reasoning in Puzzlehunts

Abstract
Puzzlehunts are a genre of complex, multi-step puzzles lacking well-defined problem definitions. In contrast to conventional reasoning benchmarks consisting of tasks with clear instructions, puzzlehunts require models to discover the underlying problem structure from multimodal evidence and iterative reasoning, mirroring real-world domains such as scientific discovery, exploratory data analysis, or investigative problem-solving. Despite recent progress in foundation models, their performance on such open-ended settings remains largely untested. In this paper, we introduce PuzzleWorld, a large-scale benchmark of 667 puzzlehunt-style problems designed to assess step-by-step, open-ended, and creative multimodal reasoning. Each puzzle is annotated with the final solution, detailed reasoning traces, and cognitive skill labels, enabling holistic benchmarking and fine-grained diagnostic analysis. Most state-of-the-art models achieve only 1-2% final answer accuracy, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. To demonstrate the value of our reasoning annotations, we show that fine-tuning a small model on reasoning traces improves stepwise reasoning from 4% to 11%, while training on final answers alone degrades performance to near zero. Our error analysis reveals that current models exhibit myopic reasoning, are bottlenecked by the limitations of language-based inference, and lack sketching capabilities crucial for visual and spatial reasoning. We release PuzzleWorld at https://github.com/MIT-MI/PuzzleWorld to support future work on building more general, open-ended, and creative reasoning systems.
1 Introduction
Figure 1: Overview of PuzzleWorld: PuzzleWorld is a dataset of complex puzzles that lack explicit instructions, requiring solvers to deduce the final answer from nuanced, multimodal cues from the puzzle content as well as external domain-specific knowledge. The raw puzzles and solutions are sourced from PuzzledPint, and the solutions, which are PNG images, are transcribed into a sequence of reasoning steps by human annotators. These annotations enable us to measure the accuracy of the final answer and the step-by-step progress made towards the solution. Best-viewed zoomed in and in color, high-resolution puzzles are in Appendix B.
Puzzles are designed to challenge a solver’s reasoning ability, creativity, and problem-solving skills. While some puzzles follow rigid formats—such as crosswords, Sudoku, or logic grids—others, like puzzlehunts, are intentionally open-ended. In puzzlehunts, solvers must first infer the nature of the task from ambiguous clues embedded in textual, visual, or cultural elements before identifying and systematically executing a multi-step solution path. These puzzles span diverse modalities and topics, including wordplay, logic, programming, visual reasoning, music, and pop culture. Often taking hours or days to solve, puzzlehunts are a staple at universities, hiring events, and community competitions, and have grown into a global phenomenon through both in-person and virtual formats.
Beyond their entertainment value, puzzlehunts provide a high-ceiling challenge for evaluating human—and increasingly, artificial—reasoning. They demand compositional thinking, lateral reasoning, trial-and-error, and backtracking dead ends. Unlike most AI benchmarks that present well-specified tasks with clear instructions, puzzlehunts require solvers to discover both what the problem is and how to solve it. This makes them uniquely well-suited for evaluating general-purpose reasoning systems under conditions that more closely resemble real-world open-ended scenarios, such as scientific discovery, investigative analysis, or exploratory problem-solving.
Recent advances in language and multimodal reasoning [22] have enabled significant progress in step-by-step problem-solving [40, 44], transparent reasoning [10, 24], and enhanced human-AI collaboration [42, 6]. However, existing reasoning benchmarks largely focus on narrow domains with fully specified tasks, such as math [23] or coding [18]. These settings do not test a model’s ability to form hypotheses, adapt to implicit structures, or reason creatively across modalities. We argue that progress toward generalist AI requires new benchmarks that challenge models to operate in less structured, discovery-driven environments that demand flexible and holistic reasoning [26].
To this end, we introduce PuzzleWorld, a benchmark of 667 real-world puzzlehunt problems curated from Puzzled Pint [29], a monthly puzzlehunt event series with content released under a Creative Commons license. These puzzles offer an open-ended, compositional challenge beyond prior benchmarks focused on instruction-following or task completion, and will grow with new puzzle releases over time. We collected fine-grained annotations of each puzzle for its final answer, the input modalities involved, the cognitive reasoning skills it exercises, and a manually curated step-by-step solution trace. These rich annotations support diagnostic analysis, model training, and detailed evaluation of intermediate reasoning capabilities.
PuzzleWorld enables us to systematically study the multimodal and multi-step reasoning capabilities of today’s best foundation models. Most state-of-the-art models achieve only 1-2% final answer accuracy, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. We additionally find that detailed annotations are important, as fine-tuning a model on annotated reasoning traces doubles a small model’s performance, while learning directly from final solutions can hurt the model and drive the accuracy towards zero. We also conduct detailed error analysis on models’ performance on PuzzleWorld, yielding tangible directions for future work in improving multimodal open-ended reasoning in AI. Together, these elements position PuzzleWorld as a rigorous resource for evaluating and improving general-purpose multimodal reasoning in AI systems. PuzzleWorld is publicly available at https://github.com/MIT-MI/PuzzleWorld. In the long run, we believe PuzzleWorld can catalyze more general and adaptable AI for mathematical and logical reasoning, open-ended scientific discovery, and autonomous assistive agents.
2 Related work
Large Language Model (LLM) Reasoning.
LLMs have demonstrated remarkable emergent capabilities, often matching or even surpassing human performance across a wide range of tasks [32]. Notably, models such as GPT-4 [2] and Claude [1] have achieved strong results not only on traditional NLP benchmarks—like question answering, summarization, and translation [41, 31, 30], but also in more complex domains such as mathematical reasoning, programming, and logical deduction [3, 18, 19]. These abilities suggest that LLMs are beginning to exhibit general-purpose reasoning skills, making them increasingly relevant to both academic research and practical applications.
This growing interest in reasoning evaluation has coincided with the rapid expansion of the LLM ecosystem, which includes both open and closed-source models. Open-source models such as LLaMA [36], Mistral [17], and Qwen [4] have become increasingly capable and accessible, while specialized variants like DeepSeek-R1 [14] are explicitly designed to improve reasoning by integrating reinforcement learning and other post-training strategies. Meanwhile, closed-source models such as GPT-o3 [28], Claude 3 Sonnet [1], and Gemini Thinking [12] have pushed the boundaries of real-world performance. However, despite these impressive capabilities, understanding the full extent and limitations of LLM reasoning remains a crucial open question, underscoring the need for benchmarks that rigorously assess their capability for flexible, holistic reasoning [26, 5].
Reasoning Benchmarks.
Numerous reasoning benchmarks have been proposed to evaluate various cognitive skills, including visual mathematical reasoning [23], spatial understanding [38], analogical reasoning [46], and social reasoning [20, 25]. However, few have addressed abstract, open-ended problems that demand holistic reasoning. HEMM [21], SciBench [39], MMMU [47], MMMU-Pro [48], MMT-Bench [45], and OlympiadBench [15] test multimodal reasoning across various disciplines in academic and real-world contexts. While these tasks are broad and challenging, they typically involve well-defined questions and are grounded in familiar domains that closely resemble the training distributions of large models. As such, they primarily assess in-distribution reasoning rather than creativity or adaptability. ARC-AGI [9] tests the ability to reason and adapt to new situations through abstract visual pattern recognition tasks that require minimal prior knowledge, yet it lacks the open-ended, exploratory nature of real-world problem solving. In contrast, PuzzleWorld  targets open-ended reasoning through puzzlehunts that lack explicit instructions or solution paths. Solving these tasks requires creatively piecing together subtle hints, often across many modalities, into coherent multi-step reasoning chains.
Figure 2: Overview of samples from PuzzleWorld Left: To gain a deeper understanding of model performance on PuzzleWorld, each puzzle is annotated with the input modalities of the puzzle content, the reasoning skills required to solve the puzzle, and step-by-step reasoning steps. Right: Example modality and reasoning skill annotations on three puzzles. High-resolution puzzle images are in Appendix B.
Puzzle Benchmarks.
A growing line of work has explored the use of puzzles to test the reasoning capabilities of AI systems. PuzzleVQA [8] consists of 2k puzzles that require abstracting patterns based on colors, numbers, shapes, and other concepts from visual puzzles to answer multiple-choice questions. AlgoVQA [13] is a benchmark of visual puzzles that require algorithmic reasoning. PUZZLES [11] tests the ability of RL agents to perform algorithmic reasoning on a set of 40 puzzles. While valuable for evaluating specific skills, these benchmarks focus on narrow domains with constrained task formats, and modern models generally perform quite well on these benchmarks [8, 27, 47]. On the other hand, the unstructured nature of the puzzlehunt problems in PuzzleWorld requires models to interpret ambiguous cues, explore creative solution strategies, and integrate information across diverse modalities and knowledge areas. These tasks often demand lateral thinking, symbolic abstraction, and visual-spatial reasoning - capabilities that go well beyond conventional pattern recognition or formal logic. The closest to our benchmark is EnigmaEval [37], which also evaluates AI’s reasoning capabilities on puzzles. However, EnigmaEval is a closed-source evaluation-only dataset and does not include manually annotated step-by-step solutions. The open-source puzzles and rich annotations in PuzzleWorld support fine-grained analysis of intermediate reasoning processes and common failure modes, facilitating the development and evaluation of more robust, general-purpose reasoning models.
3 Taxonomizing Multimodal Reasoning in Puzzlehunts
To understand how solving puzzlehunts engages reasoning capabilities evaluated separately in benchmarks like MMMU [48] and MathVista [23], we analyze puzzle solutions and classify them along two orthogonal dimensions: input modality and reasoning mechanism. This dual-axis taxonomy provides a comprehensive evaluation framework that captures both the form in which information is presented and the cognitive strategies required for reasoning.
3.1 Puzzle input modalities
To characterize the diverse nature of puzzles in our dataset, we categorize input content into three modalities: Text, which encompasses purely textual information such as written instructions, narratives, clues, or word puzzles, testing the model’s ability to comprehend and extract relevant linguistic information; Visual, which includes unstructured visual elements like images, icons, diagrams, and typography, challenging the model’s capacity to interpret visual semantics and patterns; and Structured, which refers to systematically organized visual information—such as tables, graphs, grids, matrices, and charts—requiring the model to understand relationships between data elements within defined structures. Table 1 shows the distribution of puzzles across modality and difficulty.
3.2 Puzzle reasoning mechanisms
Table 1: Count of puzzles across modalities and difficulties. Across all modalities, the distribution of difficulties is similar.
Easy Medium Hard
Text 131 322 151
Visual 90 226 111
Structured 59 181 108
We identify six core cognitive abilities essential for effective puzzle-solving in PuzzleWorld. These include logic, which covers deductive and inferential reasoning such as pattern recognition and causal inference; wordplay, involving flexible linguistic interpretation through puns, anagrams, and homophones; spatial reasoning, which tests an AI’s ability to mentally manipulate objects and navigate geometric structures; and cryptic decoding, which requires recognizing and applying transformations like ciphers and hidden encodings. In addition, knowledge-based reasoning leverages domain-specific facts from areas such as science or history, while commonsense reasoning draws on implicit real-world expectations.
This taxonomic approach enables quantitative analysis of puzzle complexity and facilitates targeted evaluation of AI reasoning capabilities across different cognitive dimensions. By mapping specific puzzles and reasoning tasks to combinations of modalities and mechanisms, we can identify areas of strength and weakness in AI systems, track progress over time, and guide future development efforts toward more balanced reasoning capabilities.
4 Creating PuzzleWorld
Statistic Value
Total # of puzzles 667
Avg. # of Reasoning Steps 5.4
Percent # of Visual Reasoning Steps 12.3%
Avg. Word Count per Reasoning Step 22.5
Correlation between Difficulty and # of Reasoning Steps 0.24
Figure 3: Dataset construction procedure and statistics: Left: First, we source raw puzzles and solutions from PuzzledPint. As the PuzzledPint solutions are in PDF format and are often not correctly parsed by OCR (for example, some solutions consist of annotated figures rather than a textual description), the metadata and reasoning steps for each puzzle are human-annotated. We use GPT-4 to automatically flag puzzles that have unclear reasoning steps and incorrect flavor texts. Finally, two human verifiers perform a manual data cleaning on the flagged puzzles to ensure a consistent annotation format. Right: We summarize the statistics of our dataset. In particular, the average number of reasoning steps is high, and the steps are relatively complex, as shown by the high average word count.
4.1 Data collection and pre-processing
We collected our puzzle corpus from Puzzled Pint [29], an organization that publishes puzzles under the Creative Commons license (CC BY-NC-SA Intl. 4.0). Their repository contains monthly puzzle sets designed for collaborative solving in public venues, covering a diverse range of puzzle types and difficulties. This allowed us to obtain more than 700 raw puzzles spanning from 2010 to 2025.
Each puzzle in our dataset consists of its original PDF containing the puzzle content, a canonical single-phrase answer, and a corresponding solution document (either a text description or an annotated PDF explaining the solution process). Unlike Wang et al. [37], we deliberately preserved the original puzzle format rather than transcribing content into separate text and images. This decision was motivated by the observation that puzzle layouts often contain crucial spatial relationships and design elements integral to the solving process. Furthermore, Wang et al. [37] demonstrated that the best foundation models are not primarily constrained by OCR capabilities. Instead, we devote our manual annotation to construct fine-grained annotations of reasoning steps extracted from existing solution documents, ensuring that the annotations accurately capture the intended solution pathways while maintaining the integrity of the original puzzle presentation.
4.2 Data annotation
To facilitate the development of reasoning capabilities in AI systems, we designed a comprehensive annotation structure for our puzzle dataset. Each puzzle in the corpus is represented by a standardized set of metadata and accompanying visual assets, enabling systematic analysis of reasoning patterns and solution pathways. To prevent ambiguity, we discard puzzles that have incomplete solutions, multiple ground truth answers, or require physical activity to solve the puzzles. This leaves us with 667 annotated puzzles.
Figure 4: Illustration of metadata schema: All puzzles are annotated with their accompanying metadata, which includes the puzzle title, flavor text, difficulty, final answer, reasoning steps, input modalities, reasoning skills, and the PuzzledPint link to the puzzle.
4.2.1 Metadata schema
Each puzzle is annotated using a structured JSON schema comprising several key fields: a title that serves as a descriptive identifier; optional flavor text that provides narrative context; a difficulty label (easy, medium, or hard) indicating the puzzle’s complexity; a solution representing the canonical answer; a reasoning field outlining an ordered sequence of steps leading to the solution; a modality tag specifying the types of input involved; a list of skills capturing the cognitive abilities required for solving; and a source field attributing the data origin. Figure 4 illustrates an annotation schema example.
4.2.2 Reasoning annotation
A key contribution of our annotation methodology is the decomposition of puzzle-solving into discrete reasoning steps. Each step is formalized as a tuple $\langle e,f\rangle$ where $e$ represents the textual explanation and $f$ denotes an optional figure illustrating the reasoning process. To ensure consistency across annotations, we loosely require each step to begin with an atomic cognitive operation, such as pattern discovery or sketching, followed by the intermediate outcome of that operation. This structured annotation enables fine-grained analysis of a model’s reasoning trajectory.
4.3 Data verification
To ensure annotation quality and integrity, we implemented a rigorous two-stage verification protocol. First, we used GPT-4o to automatically flag each puzzle annotation for correctness and reasoning coherence. This automated screening specifically identified reasoning steps exhibiting ambiguity or logical discontinuities that might impede systematic analysis, which has flagged 12.11% of the puzzle dataset. Subsequently, two human verifiers independently reviewed all flagged annotations, applying corrections where necessary. This verification process resulted in modifications to 10.93% of the initially annotated puzzles.
As an additional quality assurance measure, we conducted manual verification of a randomly selected subset comprising 5% of the total dataset. This evaluation revealed that 96.5% of the verified annotations are marked as correct by the verifiers, demonstrating the high reliability of our annotation methodology.
4.4 Dataset statistics
We summarize key statistics in Figure 3 (right). The average number of reasoning steps is above 5, and the average word count per reasoning step is above 20, demonstrating the complexity of the reasoning traces. Additionally, 12.3% of the steps have a visual intermediate output, highlighting the importance of sketching and spatial reasoning to solve puzzles. The correlation between puzzle difficulty and # of reasoning steps is 0.24. While we expect difficulty and # of reasoning steps to be positively correlated, the magnitude of the correlation is relatively low, as the difficulty of the puzzles also stems from their open-ended nature. Figure 5 shows the distribution of puzzles by modalities, reasoning skills, number of reasoning steps, and difficulty.
Figure 5: PuzzleWorld dataset statistics. Distributions of modalities and reasoning skills are balanced across different puzzles. While the majority of puzzles are classified as medium difficulty, there are a significant number of easy and hard puzzles. The number of reasoning steps follows a long-tail distribution, with many puzzle solutions requiring more than 5 steps and some hard puzzles requiring up to 30 steps of reasoning.
4.5 Checking for dataset contamination
To assess the possibility of data contamination, we test whether GPT-o3 [28] has memorized any of the puzzles in our dataset. Specifically, inspired by prior work [33, 7], we prompt the model to reconstruct the flavor text for 40 randomly sampled puzzles out of the 84 that were answered correctly. We then use GPT-4o [2] to automatically evaluate the similarity between the reconstructed and original flavor texts. We find a reconstruction accuracy of 0%, suggesting little to no evidence of data leakage. Furthermore, since Puzzled Pint [29] publishes new puzzles on a monthly basis, our dataset can be continuously updated to mitigate the risk of model overfitting on released content.
5 Experiments
In this section, we evaluate frontier closed and open-source multimodal LLMs on the PuzzleWorld dataset. We detail the evaluation setup, present quantitative results, and conduct qualitative error analysis to understand model behavior in open-ended, multimodal puzzle reasoning.
5.1 Experimental setup
We evaluate state-of-the-art closed-source and open-source multimodal large language models on PuzzleWorld. We select GPT-o3 and GPT-4o as closed-source candidates, along with open-source models Qwen QVQ[35], InternVL3[51], and Kimi VL A3B[34]. Following the protocol in Wang et al. [37], we employ a standardized prompting strategy: the model is first introduced to the puzzle-solving task with general hints, followed by the puzzle content, including images and transcribed flavor text. See Appendix C for the full prompt.
5.1.1 Automatic evaluation metrics
Beyond the models’ final answer accuracy, we additionally evaluate the models’ solution stepwise accuracy by comparing their solution reasoning with the annotated ground truth reasoning steps. Since puzzles can have multiple solution pathways, we define the stepwise accuracy score of a candidate solution to be the last annotated reasoning step it successfully identified and executed out of all the reasoning steps.
To automatically evaluate candidate responses from multiple models, we implement an LLM-as-a-judge [49] with GPT-4o to determine the score of each candidate solution. The LLM judge takes in the grading prompt, the candidate response, and the reference reasoning steps from our dataset. For each reasoning step in the reference solution, the LLM judge then determines if the step is met by the candidate response. To evaluate the reliability of the LLM judge, we compared its stepwise accuracy scores on 20 randomly selected puzzles against human-annotated evaluations. The LLM judge achieved a Pearson correlation of $r=0.829$ ($p=6.3\times 10^{-6}$) and a mean absolute error (MAE) of $0.083$ with respect to human scores, indicating strong alignment with human judgment.
5.2 Results
5.2.1 Overall performance of frontier models
Table 2: Model performance. Accuracy (Acc) and stepwise accuracy (Step) are reported overall and per modality: text, visual, and structured. Today’s state-of-the-art models struggle significantly on PuzzleWorld- most achieve only 1-2% final answer accuracy, with the best model (GPT-o3, an expensive and slow reasoning model) solving only 14% of puzzles and reaching 40% stepwise accuracy.
Overall Text Visual Structured
Model Acc Step Acc Step Acc Step Acc Step
QVQ-72B-Preview 1.36 30.23 1.33 29.25 0.63 27.96 1.18 32.40
InternVL3-78B 0.89 15.49 0.83 14.80 0.47 14.48 1.15 17.97
Kimi VL A3B 1.33 19.10 1.16 17.91 0.94 18.84 1.72 21.41
GPT-o3 14.22 39.81 15.16 39.92 8.96 33.38 13.53 41.28
GPT-4o 1.83 22.09 1.92 20.00 0.73 20.20 2.77 28.09
We report the performance of both open-source and closed-source models in Table 2. All models exhibit extremely low final answer accuracy on PuzzleWorld, with most achieving close to 1-2%. GPT-o3 attains the highest overall accuracy at 14.22%, while the best-performing open-source model, QVQ-72B-Preview, reaches just 1.36%. Although these uniformly low scores underscore the difficulty of our benchmark, they offer limited insight into the models’ reasoning capabilities.
To address this, we introduce stepwise evaluation metrics that provide a more nuanced view of intermediate reasoning performance. These metrics reveal that models with poor final answer accuracy, such as InternVL3, can still demonstrate good intermediate reasoning, achieving up to 15.49% stepwise accuracy. Similarly, while QVQ-72B-Preview lags behind GPT-4o in final answer accuracy, it outperforms it in stepwise accuracy (30.2%), reflecting more coherent reasoning trajectories despite not reaching the correct final output. These two metrics enable PuzzleWorld to remain highly challenging while offering detailed diagnostics for model evaluation and development.
In terms of input modalities, models generally perform best on text-based puzzles, with significantly lower accuracy on puzzles involving unstructured visual inputs. Interestingly, some models’ performance improves on structured visual puzzles, such as crosswords and logic grids, where the spatial format is regular and constrained. For example, GPT-o3 performs reliably on simple crosswords that have clearly defined conventions. In contrast, puzzles involving free-form or unconventional visual layouts remain particularly difficult, with models often achieving less than half their text puzzle accuracy on these inputs. These trends highlight persistent weaknesses in visual grounding and spatial reasoning across current models.
Table 3: Fine-tuning model accuracy on PuzzleWorld. Stepwise accuracy improves significantly when fine-tuning Intern-VL3 on reasoning traces, while final answer accuracy remains unchanged.
Model Acc. Step.
Base 0.76% 4.78%
Fine-tuned (Answer-only) 0.00% 2.96%
Fine-tuned (Reasoning) 0.76% 11.00%
5.3 Improving reasoning on PuzzleWorld with fine-tuning
To explore whether PuzzleWorld can support model improvement, we fine-tuned a smaller 8B Intern-VL3 model on 80% of the dataset using annotated reasoning traces, and evaluated performance on the 20% held-out test set. As a control, we fine-tuned the same model using only the final answers, without access to intermediate reasoning traces. Full fine-tuning details are provided in Appendix C.2.
Our results highlight the value of PuzzleWorld’s reasoning annotations. Fine-tuning on full reasoning traces more than doubles the model’s stepwise accuracy—from 4.78% (base model) to 11.00%. In contrast, fine-tuning on final answers alone significantly impairs performance, reducing stepwise accuracy to 2.96% and driving final answer accuracy to zero. This suggests that answer-only training can overfit to the answer only and disrupt coherent reasoning behavior.
Despite these improvements, the final answer accuracy remains unchanged at 0.76%, underscoring both the difficulty of our benchmark and the limitations of standard fine-tuning approaches in addressing the reasoning challenges in PuzzleWorld. These findings demonstrate the importance of our detailed reasoning annotations and more advanced training paradigms. We now present a detailed error analysis to identify models’ specific failure modes on PuzzleWorld that can inform future directions for multimodal reasoning.
5.4 Detailed error analysis
Figure 6: Example puzzle errors. Left: (myopic reasoning) The model outputs an incorrect plan to solve the puzzle and is unable to successfully backtrack when it hits a dead end. Middle: (language bottleneck/lack of visual understanding) The model misinterprets the visual contents of the puzzle due to inherent limitations in language. Right: (sketching errors) While the model may produce a plausible plan, it fails to execute the right visual sketching steps to obtain correct intermediate outputs. High-resolution puzzle images are in Appendix B.
We highlight the main sources of errors by the best reasoning multimodal LLMs on PuzzleWorld, focusing on GPT-o3, which achieved best results (14% final answer accuracy). See Figure 6 for example errors from each category.
Figure 7: Stepwise accuracy distribution of GPT-o3. GPT-o3 receives stepwise accuracy of 0 for most puzzles, indicating the model frequently fails to identify even the first step of the correct reasoning trace. This highlights GPT-o3’s tendency toward myopic reasoning and its inability to recover or backtrack once committed to an incorrect path.
Strong models exhibit myopic reasoning.
Despite strong performance on conventional reasoning benchmarks, frontier multimodal models often exhibit myopic commitment in their chain-of-thought behavior. Rather than flexibly exploring alternatives or revisiting prior steps, models tend to fixate on an early hypothesis–often triggered by surface-level cues–resulting in reasoning trajectories that are locally coherent but globally misaligned with the puzzle content. For example, in Figure 6, solving the puzzle requires interpreting musical notes using a mix of binary, Morse code, and flag semaphores. Instead, GPT-o3 identifies a Morse code reference early on and rigidly adheres to it–even as contradictions arise–demonstrating a lack of backtracking and global verification.
To further examine this behavior, we analyze the stepwise accuracy distribution of GPT-o3 (Figure 7). We find that, on most puzzles, the model receive a score of 0, meaning the model often fails to correctly identify even the first step of the annotated solution. Once committed to an incorrect path, the model rarely recovers, highlighting the brittle nature of its reasoning and its lack of mechanisms for self-correction or consistency checking.
Multimodal models are bottlenecked by the limitations of language.
Modern multimodal models rely heavily on language-based strategies, such as long-form chain-of-thought and code generation, for reasoning. However, we find that this dependence becomes a bottleneck in puzzles with complex visual structure. In Figure 6, the puzzle is composed of four interlocking, ring-shaped loops arranged in a clover-like pattern. This layout is intuitive to navigate visually, as the spatial flow naturally guides solvers through each ring. However, the same structure is difficult to represent linearly in text.
Figure 8: Example failure case where GPT-o3 fails to convert a complex structured puzzle into text.
While GPT-o3 correctly solves the word clues within each loop, it fails to capture the global layout when converting the puzzle into text, as shown in Figure 8. This ultimately leads the model to derive an incorrect final answer. This example highlights a broader limitation: when faced with highly structured multimodal inputs, models that default to textual reasoning often lose critical spatial information, revealing the inherent mismatch between visual intuition and language-centric inference.
Multimodal reasoning needs sketching.
While frontier models have made notable progress in logical deduction and arithmetic reasoning, they consistently underperform on spatial tasks that require sketching, drawing, and manipulating visual structure, such as decoding based on spatial arrangements or tracing paths through grids and mazes. In Figure 6, for instance, the model correctly solves the individual clues in a grid-based puzzle but fails to trace the intended path, resulting in an incorrect final answer. Humans naturally rely on sketching or mental imagery to reason through such spatial challenges, using external or internal visualizations to keep track of evolving structure. The absence of such capabilities in current models reveals a critical gap: without the ability to sketch and update a persistent visual representation during reasoning, models are prone to failure in tasks that depend on spatial coherence.
Figure 9: Reasoning skills of failed steps. We annotated the reasoning steps responsible for 30 puzzle errors with their corresponding reasoning skills.
To better understand the role of sketching in model performance, we manually analyzed 30 puzzles where GPT-o3 produced incorrect answers. For each failure case, we annotated the reasoning step responsible for the error with its corresponding reasoning skill category from the dataset. As shown in Figure 9, we found that 53.33% of these bottleneck steps involved spatial reasoning or sketching-related capabilities. This highlights a significant gap in current models’ ability to maintain and manipulate visual structure during inference. Incorporating sketch-like visual memory and spatial reasoning mechanisms [43, 16, 6] may offer a promising direction for building more robust and spatially grounded multimodal reasoning AI.
6 Conclusion
This paper presents PuzzleWorld, a large-scale benchmark of 667 puzzlehunt-style problems designed to assess step-by-step, open-ended, and creative multimodal reasoning. The diversity of puzzles and richly annotated reasoning traces enable holistic benchmarking and fine-grained diagnostic analysis. PuzzleWorld presents a unique challenge to modern LLM and multimodal model reasoning, with the best model solving only 14% of puzzles and reaching 40% stepwise accuracy. Our error analysis reveals that current models exhibit myopic reasoning, are bottlenecked by the limitations of language-based inference, and lack sketching capabilities crucial for visual and spatial reasoning. This makes PuzzleWorld uniquely well-suited for evaluating general-purpose reasoning systems under conditions that more closely resemble real-world open-ended scenarios, such as scientific discovery, exploratory data analysis, or investigative problem-solving.
Appendix A Limitations and Broader Impact
To ensure consistency and standardization across the dataset, we excluded puzzles involving underexplored or difficult-to-represent modalities such as audio, video, or interactive file-based inputs. As a result, PuzzleWorld may not fully capture the breadth of sensory and interaction-based reasoning found in some real-world, more challenging puzzlehunts. Additionally, unlike Wang et al. [37] that uses human annotators to transcribe textual and visual components separately, we preserve the puzzle content in its original image format and focus annotation efforts on intermediate reasoning traces. While this allows PuzzleWorld to provide richer annotation of the solution reasoning process, it may also introduce variability in model performance depending on the quality of their OCR capabilities. Finally, our evaluation pipeline relies on LLM-based judges to automatically assess generated reasoning traces. Although we adopt careful prompting and cross-checking, such evaluations may be subject to instability or bias, especially in edge cases involving ambiguous reasoning steps.
Our goal in releasing PuzzleWorld is to advance research in general-purpose, multimodal reasoning systems. However, we recognize that increasingly capable AI models, especially those skilled at complex reasoning, carry risks of misuse. These include the potential for externalizing or replacing human reasoning in settings where authenticity or creativity is essential, such as education, scientific authorship, or collaborative problem-solving. While our dataset does not pose direct risks on its own, we support future work that includes safeguards to mitigate misuse and encourages the responsible deployment of reasoning-capable AI systems in alignment with human values.
Appendix B PuzzleWorld details
B.1 PuzzleWorld Image Samples
We provide high-resolution images of puzzle samples used in this paper.
Appendix C Benchmarking Details
C.1 Compute Resources
All evaluations and experiments in this paper were conducted on a remote cluster equipped with two NVIDIA H200 GPUs (each with 141 GB HBM3 memory). Runtime for each model evaluation varied between 10–48 hours depending on the model size and architecture.
C.2 Supervised Finetuning
All fine-tuning experiments were conducted using the LLaMA Factory framework [50]. For the 8B InternVL3 model, we used LoRA fine-tuning with a rank of 8, a learning rate of $1\times 10^{-6}$, and trained for 3 epochs on PuzzleWorld. No additional hyperparameter tuning was performed.
C.3 Prompt for benchmarking
Below is the system prompt template for benchmarking models on PuzzleWorld, which is adapted from Wang et al. [37].
You will be presented with a puzzle to solve. The puzzle may not have specific instructions,
but you know that the answer to the puzzle is a word or short phrase (or rarely, a number).
Do not ask any questions about how to proceed, just do your best to solve the puzzle.
Here are some tips for solving puzzles of this type:
General Tips:
- Puzzles will often have multiple steps to get to the answer word. You can usually tell you
are on the right track if the intermediate answers agree with the title, flavor, or theme
of the puzzle.
- You can usually find hints in the introductory text. For example references to "in the dark"
or "sight" are often hints something is encoded with braille.
- Puzzles often incorporate acrostics: a clue where the first letter, syllable, or word of
each line, paragraph, or other recurring feature spells out a word or message.
- If you end up with a garbled "alphabet soup", then look for a clue on how to order them.
- Indexing is one of the most common puzzle mechanisms. Try indexing when you have a list of
words or phrases and a corresponding list of numbers. Count into the word or phrase by the
given number and record the letter in that position. For example: "2 Cake, 6 Pudding, 5
Shortening" gives you "ant".
- Alpha-numeric codes are also very common. If you end up with a list of numbers try replacing
the numbers with the corresponding letters like this: 1 = A, 2 = B, 3 = C... 26 = Z.
Occasionally, these types of codes will "wrap around", so dont despair if you see a
number greater than 26. Just subtract 26 and try again. In this scenario 27 (27-26 = 1) =
A, 28 (28-26 = 2) = B etc. If you try this and it doesnt work, try other numeric codes
such as ASCII.
- Often a puzzle repeats a strategy multiple times.
You will likely need to backtrack frequently, so make sure to write out your steps as you go.
If you get stuck, try to think of a new way to approach the puzzle. Try:
- Rereading the title and the flavor text. These are the most important hints about what type
of strategies, themes or cultural references might be used to solve the puzzle.
- Checking for pop culture references
- Checking for references to a song/poem/book/movie/TV show
For strings, examples of strategies you might try include:
- Alphabetizing
- Using leftover letters to spell something
- Rearranging the letters (aka anagrams or "transposing")
- Seeing if there are any acronyms
- Diagonalizing (taking the first letter of the first answer, the second letter of the second
answer, etc.)
- Looking for unusual letter frequencies
- Puns and homophones
- Shifting from letters to numbers
For numbers, try:
- Shifting from numbers to letters
- Using it as a phone number
- Treating numbers as dates
- Treating numbers as ASCII numbers
- Seeing if there are any strange sequences
- Seeing if prime numbers are involved
For images, try:
- Looking at it in a mirror
- Squinting at it from far away
- Tilting it
- Looking at it upside down
- Looking through it
- Transcribing it neatly
We additionally append the user prompt:
Your task is to solve the following puzzle. The attached images are presented in the order
they are referenced in the text.
The puzzles title is: {}
The puzzles flavor text is: {}
Write out a step-by-step solution to the puzzle. At the end of your solution, write your
answer in the following format:
Answer: <answer>
Below is the prompt for LLM judge:
Answer Equivalence Instructions:
Using the puzzle and the reference solution, grade the candidate solution as follows.
For every reasoning step of the reference solution, output True if the candidate solution both includes
the step and achieves the same intermediate result of the step, otherwise False.
Explain why the candidates solution did or did not get the reasoning step correct.
Do not add more steps than there are in the reference solution and evaluate every step
in the reference solution.
There is a exception in scoring for the last reasoning step. Identify the candidate output solution.
If the candiate output solution is the exact same as the reference solution answer of \"{puzzle_solution}\",
then output final step as true.
Appendix D Annotator Details
We employ university undergraduates to assist the human annotation process in PuzzleWorld. All annotators are compensated at a rate of \$16.00 per hour. Prior to annotation, annotators receive detailed guidelines and participate in training sessions to ensure consistency and task understanding.
D.1 Annotator Instructions
We provide the instructions given to annotators below:
# Instructions for Submitting a Puzzle
To submit a puzzle, fork this repository and create a new branch. Then, create a new folder ‘{puzzle_name}‘ in the data/puzzles folder, and place the following files in it:
- metadata.json‘: A JSON file containing the metadata of the puzzle
- content.png‘: The image of the puzzle content
[…]
For an example puzzle, see the data/puzzles/example folder. After you are done, create a pull request to merge your branch into the main repository.
Note, please replace any spaces in the puzzle name with _ when creating the new folder!
## Metadata
The metadata.json file should contain a JSON object with the following fields:
| Field Name | Type | Description |
| title | string | The title of the puzzle |
| flavor text | string | The flavor text of the puzzle, possibly empty |
| difficulty | string | The difficulty level of the puzzle (easy, medium, hard) |
| solution | string | The solution to the puzzle |
| reasoning | Step\[ \] | An ordered list of reasoning [steps](#reasoning-step) towards the solution |
| modality | string\[ \] | A list of input [modalities](#a-list-of-input-modalities) the puzzle contains |
| skills | string\[ \] | A list of [skills](#a-list-of-reasoning-skills) required to solve the puzzle |
| source | url | Thel link to the puzzle |
### Reasoning Step
The reasoning field should contain a list of Step objects, which are represented as dictionaries with the following fields:
| Field Name | Type | Description |
| explanation | string | The textual explanation of the step |
| figure | file path | (Optional) File path to a figure illustrating the step |
Each of the explanation should begin with one of the following atomic actions:
- Pattern discovery: discover patterns / insights from current information
- E.g. discovering that current laser patterns are semaphores
- Sketching: sketching on or interacting with visual elements
- E.g. traversing through a maze
- E.g. connecting the dots
- Manipulation: manipulating or arranging a sequence of elements
- E.g. sorting alphabets in order
- E.g. applying cryptic encoding / decoding
- Combining / Chaining: combining or chaining multiple pieces of observations
- E.g. matching patterns in images with text segments
- Extraction: extracting information from one pattern or observation
- E.g. extracting letters from semaphore patterns
(Note: the exact wording of action is not important as long as it resembles one of the above categories)
Each explanation step should consist of one action and the intermediate outcome of the action e.g. Identify the pattern that (...), which is (...)
### A List of Input Modalities
| Keyword | Description |
| text | Textual information |
| visual | Unstructued visual information e.g. images, icons, fonts, etc. |
| structured | Structured visual information e.g. tables, graphs, crosswords, etc.|
### A List of Reasoning Skills
| Keyword | Description |
| logic | Logic reasoning e.g. rule deduction or inferring conclusion given partial information |
| wordplay | Manipulating words based on linguistic properties e.g. anagrams, homophones, etc. |
| spatial | Spatial or visual understanding, manipulation and navigation e.g. mazes, connecting dots, etc. |
| cryptic | Encoding and decoding information e.g. ciphers, indexing, etc. |
| knowledge | Leverarging domain-specific knowledge e.g. history, science, etc. |
| commonsense | Applying common sense reasoning e.g. physical laws, social norms, etc. |
| tool_use | Searching through an external database for information unlikely in models training data, such as Google Maps |